OpenStudy (alesia):

whats the first derivative of y=sin (1+3x)^2

OpenStudy (anonymous):

To solve this you need to use the "outside-inside" rule, or the chain rule.

OpenStudy (phi):

did you get this one? if this were just y= sin(x) dy = d sin(x) dy= cos(x) dx or dy/dx = cos(x) if we let u= (1+3x)^2 y= sin(u) dy= cos(u) du or (replace u) dy= cos (1+3x)^2 d(1+3x)^2 can you finish?

OpenStudy (anonymous):

the sine isn't being squared?

OpenStudy (phi):

I am assuming the problem is \[ \frac{d}{dx}\sin( (1+3x)^2) \] if it is \[ \frac{d}{dx}\sin^2( 1+3x) \] that is a different problem